OpenStudy (anonymous):

SOMEONE HELP ME PLEASE!!!!!!!!!!!!!!!!!!: express the ratio $4 for 8 cans of soup.

OpenStudy (johnweldon1993):

well $4 for 8 cans of soup... for 8$ how many cans will we have? double the 8 right? so 16 now for $2 how many cans will we have? ...half that 8...so 4 and finally for $1...how many cans? ..half that 4...so 2 so the ratio is 1:2

OpenStudy (johnweldon1993):

in other words...for every dollar spent....you get 2 cans of soup

OpenStudy (anonymous):

I have to set it up. like something over something.

OpenStudy (johnweldon1993):

okay...$4 for 8 cans 4/8 simplifies down to 1/2 where the numerator is the number of dollars...the denominator is the number of soup cans

OpenStudy (anonymous):

ohhh, so $2 a can.

OpenStudy (anonymous):

can you help me with a few more problems?

OpenStudy (johnweldon1993):

....no...1 dollar every 2 cans...or...for every 1 can...50 cents

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

an apple crisp recipe calls for 20 pounds of apples for 30 servings. how many pounds of apples are needed for 60 servings?

OpenStudy (anonymous):

it's 40 I think but idk how to show the work.

OpenStudy (johnweldon1993):

you are correct....and just think about it...if you need 20 pounds for 30 servings...for DOUBLE the servings...you would need DOUBLE the pounds of apples...so yes 40

OpenStudy (johnweldon1993):

or to show the work 20/30 = x/60 now simplify 2/3 = x/60 what is 2/3 times 60? 60 * 2 = 120....120 / 3 = 40

OpenStudy (anonymous):

:) thanks. pia used a scale of 1 inch = 2 feet to construct a scale model of her backyard. a tree in the model has a height of 4 inches. what is the actual height of the tree?

OpenStudy (anonymous):

do you know it?

OpenStudy (johnweldon1993):

sorry for the late reply, and just in case you didnt figure this one out yet... if pia makes 1 inch = 2ft doubling the inches (2 inches) ..would double the feet right? (4ft) doubling THAT term (2 inches...to 4 inches) would double THAT feet (4 to 8 feet. so in her scale....4 inches = 8 feet :)